c) Martin, Arop and Sam and went shopping at Kabale mega super market buying similar items. Martin bought one kilogram of sugar and one tin of blue band spending 6 shillings altogether. Arop bought bought 3 kilograms of maize flour and one tin of blue band spending 17 shillings altogether. Sam bought two kilograms of sugar, one kilogram of maize flour and 3 tins of blue band spending 15 shillings altogether. i) Express their purchases as equations ii) Using matrix method, find the cost of each kilogram of sugar, maize flour and a tin of blue band.


1
Expert's answer
2022-04-28T17:39:58-0400

x = one kilogram of sugar

y = one tin of blue band

z = one kilogram of maize

"\begin{cases}
 x+y=6\\
 y+3z = 17\\
 2x +3y+z = 15
 \end{cases}"

We rewrite the system of equations in matrix form and solve it using the Gauss method

"\begin{pmatrix}
 1 & 1 & 0 \:\: |6\:\:\\
 0 & 1 & 3 \:\:| 17 \\
 2 & 3 & 1\:\:| 6\:\:
\end{pmatrix}\\"

from line 3 subtract 1 line multiplied by 2

"\begin{pmatrix}
 1 & 1 & 0 \:\: |6\:\:\\
 0 & 1 & 3 \:\:| 17 \\
 0 & 1 & 1\:\:| 3\:\:
\end{pmatrix}\\"

from line 1 we subtract line 2, multiplied by 1; from line 3 subtract line 2 multiplied by 1

"\begin{pmatrix}
 1 & 0 & -3 \: |-11\\
 0 & 1 & 3 \:|\:\: 17 \\
 0 & 0 & -2\:| -14
\end{pmatrix}\\"

Divide the 3rd row by -2

"\begin{pmatrix}
 1 & 0 & -3 \: |-11\\
 0 & 1 & 3 \:|\:\: 17 \\
 0 & 0 & 1\:|\:\:\:\: 7
\end{pmatrix}\\"

to line 1 add line 3 multiplied by 3; from line 2 subtract line 3 multiplied by 3

"\begin{pmatrix}
 1 & 0 & 0 \: |\:\:\:10\\
 0 & 1 & 0 \:| -4 \\
 0 & 0 & 1\:|\:\:\:\:\: 7
\end{pmatrix}\\"

twet

"\begin{cases}
 10\\
 -4\\
 7
 \end{cases}"

ANSWER:

x = 10

y = -4

z = 7

BUT "y" cannot be negative